# Leveraging Anthropometric Measurements to Improve Human Mesh Estimation and Ensure Consistent Body Shapes

### v1

@@ Abstract
The basic body shape of a person does not change within a single video. However, most SOTA human mesh estimation (HME) models output a slightly different body shape for each video frame, which results in inconsistent body shapes for the same person. In contrast, we leverage anthropometric measurements like tailors are already obtaining from humans for centuries. We create a model called A2B that converts such anthropometric measurements to body shape parameters of human mesh models. Moreover, we find that finetuned SOTA 3D human pose estimation (HPE) models outperform HME models regarding the precision of the estimated keypoints. We show that applying inverse kinematics (IK) to the results of such a 3D HPE model and combining the resulting body pose with the A2B body shape leads to superior and consistent human meshes for challenging datasets like ASPset or fit3D, where we can lower the MPJPE by over 30 mm compared to SOTA HME models. Further, replacing HME models estimates of the body shape parameters with A2B model results not only increases the performance of these HME models, but also leads to consistent body shapes.
Figure 1: Two qualitative examples from the ASPSet sports dataset. The result from a SOTA HME model, SMPLer-X (Cai et al., 2024), is shown on the left, the result from our model on the right, respectively. GT joints and estimated joints are color-coded. Corresponding joints are connected.
@@ 1 Introduction
Creating an accurate 3D mesh from monocular images or videos creates new opportunities in fields like 3D animation, gaming, fashion, sports, etc. In many of these application fields, videos are of main interest. While applying HME to videos, one would expect to get a mesh sequence per person with those meshes just changing their poses, but not their body shape. However, analyses of results of SOTA HME models in challenging scenarios show that the body shape of the meshes of the same person differs from frame to frame. This is because most HME models work per image and are unable to process a whole video. Worse, an analysis of currently used 3D mesh and pose datasets reveals the same inconsistencies in the provided ground truth (GT) data. For a precise body posture analysis, as it is necessary in many sports disciplines, an exact model of the athlete’s body shape is required. Therefore, many professional athletes are measured anthropometrically during performance assessments. Moreover, the body shape of an actor performing motions for 3D animations needs to be consistent as the basic body shapes does not change during performances. Changes are only due to different poses and are modeled separately besides the core shape. Thus, the changing body shapes of HME models for the same person are highly unwanted and simply wrong.
@@
Our work aims to create a single perfectly fitting basic body shape for each human and reuse it for all videos with this person. Measuring the human body has already been done for centuries to fit suits or dresses perfectly to a specific body shape. Athletes are also measured for decades for precise performance assessments. In many applications, measuring the person in action would add only a marginal overhead, but ease the postprocessing steps. This is the reason why we propose to use these measurements. However, body shape parameters of common human mesh models like SMPL-X (Pavlakos et al., 2019) are not human interpretable. Therefore, it is not possible to obtain the perfect body shape parameters by anthropometric measurements. Hence, we train a machine learning model (called A2B, anthropometric measurements to body shape) to translate those measurements into body shape parameters for HME. With this approach, measuring a person once creates the body shape that can be used for all frames in all evaluation videos.
@@
HME models are currently performing well on everyday data. However, in more challenging scenarios like sports, their performance is still inferior to finetuned SOTA 3D HPE models. 3D HPE models only predict 3D keypoints resulting in a stickfigure pose, whereas HME models output a posed mesh including the human’s surface. Due to the lack of GT meshes, HME models cannot be trained on datasets with solely 3D keypoint annotations which are easier to create than mesh annotations. The usage of synthetic data is emerging in the field, but is not applicable to challenging or specific scenarios like sports. In this paper, we propose a solution to that problem. With our A2B model and anthropometric measurements, we can now create the body shape parameters of humans needed for HME. We further apply inverse kinematics (IK) to produce the rotations that are missing in the 3D stick-figure model that is created by 3D HPE models. Together with our A2B body shape, we are able to generate human meshes that have a consistent body shape and a precise pose. Evaluations show that the performance of this composed approach is superior compared to HME models. We show qualitative examples of our model and a SOTA HME model, SMPLer-X (Cai et al., 2024), in Figure 1. Our approach is generally applicable to any HME problem. We choose sports datasets to validate our proposed approach, since the poses in sports change rapidly, which exacerbates the problem of inconsistent body shapes. Our contributions can be summarized as follows:
@@
  • We reveal inconsistencies in the GT data of ASPset (Nibali et al., 2021) and fit3D (Fieraru et al., 2021). The body shape of a single person varies mistakenly in the GT.
@@
  • We create and evaluate different models for converting between anthropometric measurements and SMPL-X body shape parameters. We call these models A2B. We are the first to intoduce a possibility to convert from anthropometric measurements to body shape parameters.
@@
  • With finetuned SOTA 2D and 3D HPE models (Xu et al., 2022; Einfalt et al., 2023), inverse kinematics (IK), anthropometric measurements, and our A2B model, we estimate accurate human meshes with a consistent body shape. We show that this approach achieves superior results to SOTA HME models, although still evaluated on the inconsistent GT.
@@ 2 Related Work
Human Mesh Estimation (HME) is an active area of research. Body models like SMPL (Loper et al., 2015) and its successor SMPL-X (Pavlakos et al., 2019) are broadly used. Their advantage is that they decouple human pose and shape. The pose parameters $\theta$ give the rotations of the joints relative to the parent joint. The shape parameters $\beta$ model the basic body shape. At first, a mesh is created with a linear mapping from $\beta$ parameters to a T-shaped pose. Next, some pose-specific shape deformations are applied and then the mesh is rotated at the joints according to the $\theta$ parameters.
@@
The first HME model that estimates SMPL-X meshes from images was introduced along SMPL-X (Pavlakos et al., 2019). This model, called SMPLify-X, detects 2D image features and then fits an SMPL-X model to these. To achieve that, they incorporate a pose prior trained on a large motion capture dataset and an interpenetration test. A more recent model for HME is Multi-HMR (Baradel et al., 2024). It predicts 2D heatmaps for person centers and based on that the human mesh with a human prediction head. OSX (Lin et al., 2023) is another HME model. It uses a component aware transformer that is composed of a global body encoder and local decoders for face and hands. SMPLer-X (Cai et al., 2024) is introduced as a generalist foundation model for HME, which is trained on a large amount of datasets. It mainly uses vision transformers. There is a multitude of other recent HME models, some focussing more on whole-body HME (Choutas et al., 2020; Feng et al., 2021; Moon et al., 2022a), others on multi-person HME - either with a two stage approach using a person detector and a single person human mesh estimator (Choi et al., 2022; Goel et al., 2023; Qiu et al., 2022), or a single stage approach estimating the meshes of all persons at once (Qiu et al., 2023; Sun et al., 2021; Zhang et al., 2021).
@@
Choutas et al. (2022) observed that existing HME models focus more on the body pose than the shape, although the shape is equally important for many applications. They propose SHAPY, a model that uses anthropometric and linguistic attributes to create accurate body shapes. Moreover, Sarkar et al. (2023) introduce SoY, which contains specific loss functions to enhance the body shape accuracy. They further propose a refinement step during test time that enhances the shape quality even more. KBody (Zioulis & O’Brien, 2023) follows a predict and optimize approach for HME. It leverages virtual joints in order to achieve that and focuses on pixel alignment to accurately capture the shape of the human. AnthroNet (Picetti et al., 2023) propose a new body model that is learned with an end-to-end trainable pipeline. It takes anthropometric measurements as an input to learn a mesh model that accurately captures shapes of humans, but this model is different from the commonly used SMPL-X model. We use the common SMPL-X body model and decouple the estimation of the shape from the estimation of the pose. This way, we ensure the consistency of the body shape over time.
@@
Inverse kinematics (IK) are common in the field of robotics. In the last years, 3D HME approaches leveraged IK to enhance their output. HybrIK (Li et al., 2021) transforms 3D joint coordinates to relative body-part rotations for 3D HME by using a twist-and-swing decomposition. HybrIK-X (Li et al., 2023b) further enhances HybrIK with expressive face and hands. Cha et al. (Cha et al., 2022) leverage IK to tackle the challenge of person-to-person occlusions in images with interacting persons. PLIKS (Shetty et al., 2023) (Pseudo-Linear Inverse Kinematic Solver) approaches HME as a model-in-the-loop optimization problem by analytically reconstructing the human model via 2D pixel-aligned vertices in an IK-like manner. NIKI (Li et al., 2023a) tries to unite robustness to occlusions and pixel-aligned accuracy in non-occlusion cases by modeling bi-directional errors. NIKI can learn from both the forward and inverse processes, whereby it emulates the analytical inverse kinematics algorithms with the twist-and-swing decomposition for better interpretability.
@@
Although HME is an active area of research, it is yet not common in computer vision for sports. Due to high velocities and a great variation of poses, sports is a challenging scenario for all kinds of human pose and shape estimation. The fit3D dataset (Fieraru et al., 2021) is a dataset which consists of videos from gym sports exercises with repetitions and is annotated with human meshes. AIFit (Fieraru et al., 2021) is a tool trained on fit3D which can reconstruct 3D human poses, reliably segment exercise repetitions, and identify the deviations between standards learned from trainers, and the execution of a trainee to give feedback to trainees. Other sports datasets only consist of 3D joint annotations, like ASPset (Nibali et al., 2021) or SportsPose (Ingwersen et al., 2023). SportsCap (Chen et al., 2021) is an approach for simultaneously capturing 3D human motions and understanding fine-grained actions from monocular challenging sports videos. Baumgartner & Klatt (2023) make use of partial sports field registration to enhance 3D HPE results for athletes running on a track.
@@ 3 Errors in 3D Human Shape Ground Truth
Each person has a specific body shape that does not change over a short time period. Therefore, the SMPL-X body model decouples the human pose encoded by $\theta$ parameters from the body shape encoded by $\beta$ parameters. Deformations to the mesh that are caused by the pose are modeled separately. Therefore, it makes sense to assign a single set of shape parameters $\beta$ to a person to describe his/her shape for a given short time period such as a recorded action. Further, there are lengths that can be calculated from 3D joints that should never change, since individual bones of humans are rigid and should not be deformed by different poses. Our approach enforces a single set of shape parameters per person and immutable bone lengths.
@@
Table 1: GT data analysis for ASPset (left) and fit3D (right): Standard deviation $\sigma$, relative standard deviation $\frac{\sigma}{\textit{avg}}$ and relative range $\frac{\max-\min}{\textit{avg}}$ of anthropometric measurements. Standard deviations are given in cm, but for the $\beta$ parameters. The values are averaged between left and right body parts and between all persons in the test set of each dataset. The $\beta$ parameter standard deviation is averaged over all $\beta$ parameters.
@@
Measurement $\sigma$ r. $\sigma$ r. range Measurement $\sigma$ r. $\sigma$ r. range
head 0.91 5.98% 57.91% head 0.69 2.60% 14.70%
hip width 1.71 9.48% 85.46% hip circ. 0.80 0.76% 5.44 %
forearm 1.99 8.37% 92.04% forearm 0.33 1.40% 7.64%
upper arm 1.72 6.29% 66.35% arm 0.76 1.53% 8.32%
lower leg 1.44 3.60% 41.36% lower leg 0.43 1.14% 11.14%
thigh 1.65 4.23 % 35.46% thigh 0.37 1.03% 9.21%
height 1.65 0.98% 6.45%
$\beta$ param. 0.65
@@
In this paper, we use ASPset (Nibali et al., 2021) and fit3D (Fieraru et al., 2021), since both datasets consist of videos with fast changing poses and 3D GT. Results for the Human3.6m (Ionescu et al., 2014) and MPI-INF-3DHP (Mehta et al., 2017) datasets are presented in the supplementary. For ASPset, we analyze bone lengths, since it has only GT annotations for 3D joints. For fit3D, GT SMPL-X $\beta$ parameters are available, hence we can analyze the $\beta$ parameters directly and further the derived anthropometric measurements. These values are the output of our deterministic B2A function: It generates a standard T-pose with the given $\beta$ parameters and computes 36 anthropometric measurements from the resulting mesh. Results of our GT analysis for a subset of the anthropometric values are shown in Table 1. We can see that the GT itself is not consistent. The deviations are larger for ASPset, but although we have GT SMPL-X meshes for fit3D, the $\beta$ parameters of a single person have a standard deviation of 0.65 on average. [^1] Averaged standard deviation means (in the whole paper) that the standard deviation is calculated per person, and the mean of the resulting standard deviations is calculated afterwards. This is a relevant flaw in the GT shape annotation, since based on the model, the GT shape should be static for each human. Nevertheless, we use the given inconsistent GT for our evaluations for comparability with related work and as we have no good means to correct them. The reader should keep this in mind. Nevertheless, we want to encourage future research in the field of 3D human pose and mesh data collection to try to eliminate these flaws in the provided GT.
4 From Anthropometric Measurements to the Body Shape
@@
Humans have been measured for centuries (Doyon et al., 2023). Tailors know exactly which measurements to take for perfectly fitting a suit or dress to the body shape of a customer. In sports, professional athletes are measured for precise performance assessments, too. Measuring a human is easy and well understood. In contrast, the parameters of the body shape for human mesh models like SMPL-X (Pavlakos et al., 2019) are not humanly interpretable. The $\beta$ parameters describe the principal components of the human body shape with typically around 10 to 16 values and are the result of a PCA executed on the human meshes of a training dataset while learning the SMPL-X model. Fixing all $\beta$ parameters despite one and looking at the results lets human observers get a notion of what this parameter might mean, but in total, the $\beta$ parameters and their interactions are not well interpretable. Therefore, we want to leverage the well established technique of measuring humans to create precise body shape parameters for the commonly used SMPL-X human mesh model. We call our approach to convert from 36 Anthropometric measurements to Body shape parameters A2B. Since there is no known relation between anthropometric measurements and $\beta$ parameters, our aim is to learn this mapping. We are the first to present a method to convert from anthropometric measurements to body shape parameters. The reverse direction, B2A, is a deterministic function of the human mesh, as the anthropometric measurements can be measured from the mesh.

### v3

@@ Abstract
The basic body shape (i.e., the body shape in T-pose) of a person does not change within a single video. However, most SOTA human mesh estimation (HME) models output a slightly different, thus inconsistent basic body shape for each video frame. Furthermore, we find that SOTA 3D human pose estimation (HPE) models outperform HME models regarding the precision of the estimated 3D keypoint positions. We solve the problem of inconsistent body shapes by leveraging anthropometric measurements like taken by tailors from humans. We create a model called A2B that converts given anthropometric measurements to basic body shape parameters of human mesh models. We obtain superior and consistent human meshes by combining the A2B model results with the keypoints of 3D HPE models using inverse kinematics. We evaluate our approach on challenging datasets like ASPset or fit3D, where we can lower the MPJPE by over 30 mm compared to SOTA HME models. Further, replacing estimates of the body shape parameters from existing HME models with A2B results not only increases the performance of these HME models, but also guarantees consistent body shapes.
@@ 1 Introduction
Creating an accurate 3D human mesh from monocular images or videos creates new opportunities in fields like 3D animation, gaming, fashion, sports, etc. In many of these application fields, videos are of main interest.
Figure 1: Two qualitative examples from the ASPSet sports dataset. The result from a SOTA HME model, SMPLer-X [3], is shown on the left, the result from our model on the right, respectively. GT joints and estimated joints are color-coded. Corresponding joints are connected.
@@
While applying HME to videos, analyses of results of SOTA HME models show that the basic body shape of the meshes of the same person differs from frame to frame.  [^1] The body shape in a given pose is usually modeled by a basic body shape (given in T-pose) plus an additional pose-dependent deformation. We call the basic body shape just body shape in the following, as the pose-specific correction is computed from the pose and does not need to be estimated. Worse, an analysis of currently used 3D mesh and pose datasets reveals the same inconsistencies in the provided ground truth (GT) data. For a precise body posture analysis, as it is necessary in many sports disciplines, an exact model of the athlete’s body shape is required. Therefore, most professional athletes are measured anthropometrically during performance assessments today. Moreover, the body shape of an actor performing motions for 3D animations needs to be consistent as the basic body shapes does not change during performances. Thus, the changing body shapes of HME models for the same person are highly unwanted and simply wrong.
@@
Our work aims to create a single perfectly fitting basic body shape for each human and reuse it for all videos with this person. Measuring the human body has already been done for centuries to fit suits or dresses perfectly to a specific body shape. In many applications, measuring the person in action beforehand would add only a marginal overhead, but improves the results dramatically. For this reason we propose to use these measurements. Body shape parameters of common human mesh models like SMPL-X [27] are not human interpretable. Therefore, it is not possible to obtain the perfect body shape parameters by anthropometric measurements. Hence, we train a machine learning model (called A2B, anthropometric measurements to body shape) to translate those measurements into body shape parameters for HME. With this approach, measuring a person once creates the body shape that can be used for all frames in all evaluation videos.
@@
HME models are performing well on everyday data. However, in more challenging scenarios like sports, their performance is inferior to finetuned SOTA 3D HPE models. 3D HPE models only predict 3D keypoints resulting in a stickfigure pose, whereas HME models output a posed mesh including the human’s surface. Due to the lack of GT meshes, HME models cannot be trained on datasets with solely 3D keypoint annotations. The usage of synthetic data is emerging in the field, but is not applicable to challenging or specific scenarios like sports. In this paper, we propose a solution to that problem. With our A2B model and anthropometric measurements, we can now create the body shape parameters of humans needed for HME. We further apply inverse kinematics (IK) to produce the rotations that are missing in the 3D stick-figure model that is created by 3D HPE models. Together with our A2B body shape, we are able to generate human meshes that have a consistent body shape and a precise pose. The main goal of this paper is to estimate the best possible human mesh with a consistent body shape by adding the marginal overhead of measuring humans. Since our main focus is sports, this overhead is negligible, as professional athletes are commonly measured anyways. We show qualitative examples of our model and a SOTA HME model, SMPLer-X [3], in Figure 1. Our approach is generally applicable to any HME problem. We choose sports datasets to validate our proposed approach, since the poses in sports are challenging for existent HME models and athletes are measured. Their performance is currently not good enough to use them in performance assessments of athletes, which we try to change. Our contributions can be summarized as follows:
@@
  • We reveal inconsistencies in the GT data of ASPset [25] and fit3D [12]. The body shape of a single person varies mistakenly in the GT.
@@
  • We create and evaluate different models to convert between anthropometric measurements and SMPL-X body shape parameters. We call these models A2B. We are the first to intoduce a possibility to convert from those measurements to body shape parameters.
@@
  • With finetuned SOTA 2D and 3D HPE models [35, 10], IK, anthropometric measurements, and our A2B model, we estimate accurate human meshes with a consistent body shape. We show that this approach achieves superior results to SOTA HME models, although still evaluated on the inconsistent GT.
@@ 2 Related Work
Human Mesh Estimation (HME) is an active area of research. Body models like SMPL [20] and its successor SMPL-X [27] are broadly used. Their advantage is that they decouple human pose and shape. The pose parameters $\theta$ give the rotations of the joints relative to the parent joint. The shape parameters $\beta$ model the basic body shape. At first, a mesh is created with a linear mapping from $\beta$ parameters to a T-shaped pose. Next, some pose-specific shape deformations are applied and then the mesh is rotated at the joints according to the $\theta$ parameters.
@@
The first HME model that estimates SMPL-X meshes from images was introduced along SMPL-X [27]. This model, called SMPLify-X, detects 2D image features and then fits an SMPL-X model to these. To achieve that, they incorporate a pose prior trained on a large motion capture dataset and an interpenetration test. A more recent model for HME is Multi-HMR [1]. It predicts 2D heatmaps for person centers and based on that the human mesh with a human prediction head. OSX [19] is another HME model. It uses a component aware transformer that is composed of a global body encoder and local decoders for face and hands. SMPLer-X [3] is introduced as a generalist foundation model for HME, which is trained on a large amount of datasets. It mainly uses vision transformers. There is a multitude of other recent HME models, some focussing more on whole-body HME [7, 11, 23], others on multi-person HME - either with a two stage approach using a person detector and a single person human mesh estimator [6, 13, 29], or a single stage approach estimating the meshes of all persons at once [30, 34, 36].
@@
Choutas et al. [8] observed that existing HME models focus more on the body pose than the shape, although the shape is equally important for many applications. They propose SHAPY, a model that uses anthropometric and linguistic attributes to create accurate body shapes. Moreover, Sarkar et al. [32] introduce SoY, which contains specific loss functions to enhance the body shape accuracy. AnthroNet [28] propose a new body model that is learned with an end-to-end trainable pipeline. It takes anthropometric measurements as an input to learn a mesh model that accurately captures shapes of humans, but this model is different from the commonly used SMPL-X model. We use the common SMPL-X body model and decouple the estimation of the shape from the estimation of the pose. This way, we ensure the consistency of the body shape over time.
@@
In the last years, 3D HME approaches leveraged Inverse kinematics (IK) to enhance their results. HybrIK [17] transforms 3D joint coordinates to relative body-part rotations for 3D HME by using a twist-and-swing decomposition. HybrIK-X [18] further enhances HybrIK with expressive face and hands. Cha et al. [4] leverage IK to tackle the challenge of person-to-person occlusions in images with interacting persons. PLIKS [33] (Pseudo-Linear Inverse Kinematic Solver) approaches HME as a model-in-the-loop optimization problem by analytically reconstructing the human model via 2D pixel-aligned vertices in an IK-like manner.
@@
Although HME is an active area of research, it is yet not common in computer vision for sports. Due to high velocities and a great variation of poses, sports is a challenging scenario for all kinds of human pose and shape estimation. The fit3D dataset [12] is a dataset which consists of videos from gym sports exercises with repetitions and is annotated with human meshes. AIFit [12] is a tool trained on fit3D which can reconstruct 3D human poses, reliably segment exercise repetitions, and identify the deviations between standards learned from trainers, and the execution of a trainee. Other sports datasets only consist of 3D joint annotations, like ASPset [25] or SportsPose [15]. SportsCap [5] is an approach for simultaneously capturing 3D human motions and understanding fine-grained actions from monocular challenging sports videos.
@@ 3 Errors in 3D Human Shape Ground Truth
Each person has a specific basic body shape that does not change over a short time period. Therefore, the SMPL-X body model decouples the human pose encoded by $\theta$ parameters from the basic body shape encoded by $\beta$ parameters. Deformations to the basic body shape that are caused by the current pose are modeled separately. Therefore, it makes sense to assign a single set of shape parameters $\beta$ to a person for a given short time period such as a recorded action to describe his/her shape. Further, there are lengths that can be calculated from 3D joints that should never change, since individual bones of humans are rigid and should not be deformed by different poses. Our approach enforces a single set of shape parameters per person and immutable bone lengths.
@@
Measure $\sigma$ r. $\sigma$ r. range Measure $\sigma$ r. $\sigma$ r. range
head 0.91 5.98% 57.91% head 0.73 2.73% 17.52%
hip width 1.71 9.48% 85.46% hip circ. 0.87 0.84% 8.17 %
forearm 1.99 8.37% 92.04% forearm 0.34 1.40% 9.24%
upper arm 1.72 6.29% 66.35% arm 0.76 1.51% 9.42%
lower leg 1.44 3.60% 41.36% lower leg 0.52 1.31% 13.80%
thigh 1.65 4.23 % 35.46% thigh 0.43 1.17% 11.74%
height 1.60 0.94% 8.69%
$\beta$ param. 0.64
Table 1: GT data analysis for ASPset (left) and fit3D (right): Standard deviation $\sigma$, relative standard deviation $\frac{\sigma}{\textit{avg}}$ and relative range $\frac{\max-\min}{\textit{avg}}$ of anthropometric measurements. Standard deviations are given in cm, but not for the $\beta$ parameters. The values are averaged between left and right body parts and between all persons used for evaluations in Section 5. The $\beta$ parameter standard deviation is averaged over all $\beta$ parameters.
@@
In this paper, we use ASPset [25] and fit3D [12], since both datasets consist of videos with fast changing poses and 3D GT. Results for the Human3.6m [16] and MPI-INF-3DHP [22] datasets are presented in the supplementary. For ASPset, we analyze bone lengths, since it has only GT annotations for 3D joints. For fit3D, GT SMPL-X $\beta$ parameters are available, hence we can analyze the $\beta$ parameters directly and further the derived anthropometric measurements. These values are the output of our deterministic B2A function: It generates a standard T-pose with the given $\beta$ parameters and computes 36 anthropometric measurements from the resulting mesh. Results of our GT analysis for a subset of the anthropometric values are shown in Table 1. We can see that the GT itself is not consistent. The deviations are larger for ASPset. Although we have GT SMPL-X meshes for fit3D, every $\beta$ parameter of a single person has a standard deviation of 0.64 on average. [^2] Averaged standard deviation means (in the whole paper) that the standard deviation is calculated per person, and the mean of the resulting standard deviations is calculated afterwards. This is a relevant flaw in the GT shape annotation, since based on the model, the GT shape should be consistent for each human. Nevertheless, we use the given inconsistent GT for our evaluations for comparability with related work and as we have no good means to correct them. The reader should keep this in mind. Nevertheless, we want to encourage future research in the field of 3D human pose and mesh data collection to try to eliminate these flaws in the provided GT.
4 From Measurements to Body Shape
@@
Humans have been measured for centuries [9]. Tailors know exactly which measurements to take for perfectly fitting a suit or dress to the body shape of a customer. In sports, it is already common practice that professional athletes are measured for precise performance assessments. Measuring a human is easy and well understood. In contrast, the parameters of the body shape for human mesh models like SMPL-X [27] are not humanly interpretable. The $\beta$ parameters describe the principal components of the human body shape with typically around 10 to 16 values and are the result of a PCA executed on the human meshes of a training dataset while learning the SMPL-X model. Fixing all $\beta$ parameters despite one and looking at the results lets human observers get a notion of what this parameter might mean, but in total, the $\beta$ parameters and their interactions are not well interpretable. Therefore, we want to leverage the well established technique of measuring humans to create precise body shape parameters for the commonly used SMPL-X human mesh model. We call our approach to convert from 36 Anthropometric measurements to Body shape parameters A2B. Since there is no known relation between anthropometric measurements (AMs) and $\beta$ parameters, our aim is to learn this mapping. We are the first to present a method to convert from anthropometric measurements to body shape parameters. The reverse direction, B2A, is a deterministic function of the human mesh, as the AMs can be measured from the mesh.